# Beyond PCA: Manifold Dimension Estimation via Local Graph Structure

### v1

@@ 4.1 Setup
In this section, we evaluate the performance of our proposed estimators, QE and TLS, on a diverse collection of synthetic and real-world datasets. For the synthetic datasets, each experiment is repeated $100$ times using randomly generated samples. To assess both bias and variance, we report the mean and standard deviation of the resulting estimates. For comparison, we also include Local PCA, CA-PCA, TwoNN, and DanCo, all of which have been identified as top-performing methods in the existing literature  [^1] Our experiments were conducted using Python. Specifically, we utilized the skdim package implementations for Local PCA, TwoNN, and DanCo, while the remaining estimators, as well as the sampling and testing procedures, were implemented by us. The code is available at: https://github.com/loong-bi/manifold-dimension-estimation. .
@@
Lastly, we evaluate our estimators on three real-world datasets commonly used in manifold dimension estimation studies. The first is the ISOMAP dataset, which consists of 698 grayscale images ($64\times 64$ pixels) of a face sculpture and is believed to have an intrinsic dimension of approximately 3. The second is a subset of the MNIST dataset, comprising 6,742 grayscale images of the digit ”1” (each of size $28\times 28$), with an estimated intrinsic dimension between 8 and 11. The third is the ISOLET dataset, which contains 6,240 vectors of length 616, representing frequency-based features extracted from audio recordings of spoken English letters, and is believed to have an intrinsic dimension between 16 and 22. The corresponding results are presented in Tables 9, 10, and 11222The results reported for DanCo, shown in brackets, are taken from Ceruti et al. (2014) and Campadelli et al. (2015), which include implementation-specific adjustments designed to stabilize performance. A direct application of DanCo using the skdim library does not reproduce these results; see Bi and Lafaye de Micheaux (2025) for further discussion..

### v2

@@ 4.1 Setup
In this section, we evaluate the performance of our proposed estimators, QE and TLS, on a diverse collection of synthetic and real-world datasets. For the synthetic datasets, each experiment is repeated $100$ times using randomly generated samples. To assess both bias and variance, we report the mean and standard deviation of the resulting estimates. For comparison, we also include Local PCA, CA-PCA, TwoNN, and DanCo, all of which have been identified as top-performing methods in the existing literature  [^2] Our experiments were conducted using Python. Specifically, we utilized the skdim package implementations for Local PCA, TwoNN, and DanCo, while the remaining estimators, as well as the sampling and testing procedures, were implemented by ourselves. The code is available at: https://github.com/loong-bi/manifold-dimension-estimation. .
Figure 6: Effect of Neighborhood Size $K$.
@@
Lastly, we evaluate our estimators on three real-world datasets commonly used in manifold dimension estimation studies. The first is the ISOMAP dataset, which consists of 698 grayscale images ($64\times 64$ pixels) of a face sculpture and is believed to have an intrinsic dimension of approximately 3. The second is a subset of the MNIST dataset, comprising 6,742 grayscale images of the digit ”1” (each of size $28\times 28$), with an estimated intrinsic dimension between 8 and 11. The third is the ISOLET dataset, which contains 6,240 vectors of length 616, representing frequency-based features extracted from audio recordings of spoken English letters, and is believed to have an intrinsic dimension between 16 and 22. The corresponding results are presented in Tables 9, 10, and 11333The results reported for DanCo, shown in brackets, are taken from Ceruti et al. (2014) and Campadelli et al. (2015), which include implementation-specific adjustments designed to stabilize performance. A direct application of DanCo using the skdim library does not reproduce these results; see Bi and Lafaye de Micheaux (2025) for further discussion..
@@ 5 Conclusion
With a better understanding of the underlying manifold, one could make more informed choices of the functions $G_{j}$, potentially leading to estimators that achieve improved performance for specific tasks beyond the two proposed in this work. However, the absence of suitable diagnostic tools currently prevents us from doing so. It remains unclear which sample statistics would provide meaningful guidance in selecting appropriate forms of $G_{j}$. Further research on this question—both theoretical and practical—is needed to advance this direction.